OpenStudy (needshelpdfdf):

(6x3 + 18x2) − 3x2 =

OpenStudy (quickstudent):

Are the 3 and 2 exponents?

OpenStudy (needshelpdfdf):

it just says find the difference

OpenStudy (danieldjpon3):

? difference?

OpenStudy (quickstudent):

Ok so then does the original equation look like this: \[\left(\begin{matrix}6x ^{3} + 18x ^{2} \\ \end{matrix}\right) - 3x ^{2} = \]

OpenStudy (needshelpdfdf):

yes

OpenStudy (danieldjpon3):

Ohh ok that makes more sense

OpenStudy (needshelpdfdf):

lol they don't know how